OpenStudy (anonymous):

What is the length of Line segment B C? (Line segment B coordinate is (-2,-3), and Line segment C is (0,2) - 3 units - 4 units - 5 units - 6 units

OpenStudy (dbzfan836):

well, first tell me what are your thoughts on this

OpenStudy (anonymous):

Well, from what I've been learning is that I have to match the coordinates on a graph, draw a line to connect them, and then if the line is diagonal I have to make a right triangle with it, and then try to figure out what each side of the triangle measures.

OpenStudy (dbzfan836):

correct. it is measured in units. where did it become difficult?

OpenStudy (anonymous):

I don't have my notes with me, and i forgot the equation I need to go by in order to find the answer.

OpenStudy (dbzfan836):

c^2 is the longest line in a triangle

OpenStudy (anonymous):

Oh! Thank you. So I guess I have to find the whole numbers for each coordinate, match them on the legs of the triangle and put them on the equation for solving?

OpenStudy (dbzfan836):

yes, and i will stay to check your work

OpenStudy (anonymous):

Okay. Hold on

OpenStudy (anonymous):

I think the shortest leg is 2 units, but I'm having trouble figure out what the longest other leg measures.

OpenStudy (dbzfan836):

alright, what is the y of the first point, point b?

OpenStudy (anonymous):

y= -3

OpenStudy (dbzfan836):

what about the y of point C?

OpenStudy (anonymous):

y=2

OpenStudy (dbzfan836):

A: the line would be -1 units long B: the line would be 5 units long

OpenStudy (anonymous):

I'm pretty sure its B

OpenStudy (dbzfan836):

maybe, but maybe not. first thing's first, test -1^2 + 2^2

OpenStudy (dbzfan836):

what would the answer be?

OpenStudy (anonymous):

-1^2+2^2=-1+4

OpenStudy (anonymous):

I'm not quite sure the other part, but my for sure guess is 3

OpenStudy (dbzfan836):

im sorry for late news, yet i made a mistake in the formula. it is a^2 * b^2 = c^2

OpenStudy (dbzfan836):

not much of a difference

OpenStudy (anonymous):

oh that's alright. Am I still good though?

OpenStudy (dbzfan836):

it would be -1*4

OpenStudy (dbzfan836):

yes, you are fine

OpenStudy (dbzfan836):

that would mean -4

OpenStudy (dbzfan836):

also known as answer B

OpenStudy (dbzfan836):

just to make sure, lets check 5^2*2^2

OpenStudy (dbzfan836):

25*4=100

OpenStudy (dbzfan836):

so the answer is B

OpenStudy (anonymous):

Okay,

OpenStudy (dbzfan836):

to triple-check, i have a software

OpenStudy (dbzfan836):

i will check the triangle's lengthe of line BC

OpenStudy (anonymous):

M'kay

OpenStudy (dbzfan836):

i didnt use it first because i would help you much

OpenStudy (anonymous):

That's alright

OpenStudy (dbzfan836):

its about -5. hmmm

OpenStudy (anonymous):

Well, I'm aloud to round.

OpenStudy (dbzfan836):

that changes this entirely

OpenStudy (dbzfan836):

It could be C or B

OpenStudy (dbzfan836):

More likely C

OpenStudy (anonymous):

Earlier for some reason i was so sure it was C actually.

OpenStudy (dbzfan836):

So you were correct from the start

OpenStudy (anonymous):

Oh, really? Well then, I guess I'm smarter than I thought I was. lol

OpenStudy (dbzfan836):

Never underestimate the mind, but you will always overestimate your eves

OpenStudy (dbzfan836):

They fool you all the time

OpenStudy (anonymous):

I know, it happens to me more than I'd like it too. haha But I still thank you for your help.

OpenStudy (dbzfan836):

np

OpenStudy (dbzfan836):

just remember medals for everyone that helps

OpenStudy (dbzfan836):

just click best response

OpenStudy (dbzfan836):

Its Free

OpenStudy (anonymous):

Don't worry, I give everyone medals who try to help me ^_^

OpenStudy (dbzfan836):

(づ。◕‿‿◕。)づ

OpenStudy (dbzfan836):

XD

OpenStudy (dbzfan836):

(╯°□°)╯︵ ┻━┻ ┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻ ┬──┬ ノ( ゜-゜ノ) (ノಠ益ಠ)ノ彡┻━┻

OpenStudy (anonymous):

Thanks and Good Night! I don't really mean "good night", I just remember hearing that from a recorded concert on the internet. lol bye

OpenStudy (dbzfan836):

lol bye